OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (freckles):

-(a) means the opposite of a

OpenStudy (anonymous):

so D

OpenStudy (freckles):

Looks good to me

OpenStudy (anonymous):

yeah D

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

mhm np

OpenStudy (anonymous):

o wait

OpenStudy (anonymous):

can you help me with 1 more

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

ok ill redo that

OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (anonymous):

(17 + h) + (–9) = 17 + (5 + (–9))

OpenStudy (anonymous):

ughh sorry

OpenStudy (anonymous):

its fine but i cant help because it looks like that*